OpenStudy (anonymous):

3/x-2 - 4/x-3 PLEASE HELP

OpenStudy (anonymous):

since the denominators are different we must establish a common denominator like so 3(x-3)/(x-2)(x-3)-4(x-2)/(x-2)(x-3) this simplifies to... 3(x-3)-4(x-2)/(x-2)(x-3)=3x-9-4x-8/(x-2)(x-3) =-x-17/(x-2)(x-3)